In a class of 28 students, 21 have a cat and 9 have a dog. There are 7 students who have a cat and a dog. What is the probability that a student has a dog given that they do not have a cat?


1
Expert's answer
2021-11-03T16:43:31-0400
"P(Dog|\overline{Cat})=\dfrac{P(Dog\cap\overline{Cat})}{P(\overline{Cat})}"


"N(Dog\cap\overline{Cat)}=2"

"N(\overline{Cat)}=28-21=7"

"P(Dog|\overline{Cat})=\dfrac{2}{7}"